OpenStudy (anonymous):

Hi. Can someone help me with problem 6a on the first assignment. I understand how they are finding A, but I don't understand how they are finding c. Any help would be appreciated.

OpenStudy (anonymous):

The problem is: Express in the form A sin (x+ c) sinx + √3cosx

OpenStudy (anonymous):

multiply and divide by 2..so u get 2(six/2 +sqrt3 cosx/2)... now sin 30 = 1/2 and cos 30 =sqrt 3 /2 so it becomes 2sin(x+30)

OpenStudy (jkristia):

I'm not sure your explanation is correct. If I double check the result using x = pi/12 (and rad mode) then I get #1 = 1.932 and #2 = 1.414. But most likely I'm just missing something

OpenStudy (jkristia):

I looked at the solution, and it is clear how A is found, but unclear how c = pi/3 is found.

OpenStudy (anonymous):

Since r=√a squared+b squared, √1 squared+√3 squared gives r=2. Since tan=sinx/cosx, arctan=cosx/sinx, which gives π/3. Plugging these values into the given formula, we get the answer: 2sin(x+π/3). I only found this after looking it up. I appreciate everyone who helped!

OpenStudy (jkristia):

isn't cosx/sinx = cot, not arctan ? I see tan = sin / cos = 1 / sqrt(3), but I still dont see how it can be sqrt(3)/1

OpenStudy (jkristia):

I think I finally figured it out. what is y if x = 0?, y = sqrt(3), since sin(x) = 0 and sqrt(3)cos(x) = sqrt(3). Since A = 2, the value for sin(x + c) = sqrt(3) / 2 = pi/3. So c = pi/3.

OpenStudy (anonymous):

It's inverse tangent. Sorry. Another way is jkristia's way. I think both get you the same answer.

OpenStudy (anonymous):

ireneb i'm having some trouble understanding this and was wondering where you looked it up?

OpenStudy (anonymous):

I honestly cannot remember where I looked it up. I remember from my high school PreCalc that inverse tangent helps you find angle measures, and is different than just regular tangent to find the side of a triangle. It's tangent to the -1, not tangent as in o/a or a/o.

OpenStudy (jkristia):

It is correct that \[\arctan (\sqrt{3}) = \frac{\pi}{3}\] Since \[\sin(\pi/3) = \frac{\sqrt{3}}{2}, \cos(\pi/3) = \frac{1}{2}\] so\[\tan(\pi/3) = \frac{\sin(\pi/3)}{\cos(\pi/3)}=\frac{\frac{\sqrt{3}}{2}}{\frac{1}{2}}=\sqrt{3}\] So by dividing sin(x) + sqrt(3)cos(x) by cos(x), you get\[\tan(x) + \sqrt{3}\] but I still cannot make the last connection. Why is tan c = sqrt(3) ?

OpenStudy (anonymous):

tan c should be π/3, which is the given solution. If you have \[\tan^{-1} =(\sqrt{3/1})\] you should get the answer of c=π/3, or 60 degrees.

OpenStudy (jkristia):

I know it is pi/3, and I know how to solve it by setting x = 0, and I know that arctan(sqrt(3)) is pi/3, - but I still dont see why it is tan c.

OpenStudy (anonymous):

You are expressing in the form A sin(x +c). We found A to be 2, and x is not what we are needing to solve for. C is just the variable that we solved for. It just means that we are going to put in π/3 for c in the required form.

OpenStudy (anonymous):

thanks ireneb ! i will try it again and c if i can't figure it out.